### v2

# Multivalent linkers mediated ultra-sensitive bio-detection

Abstract
In biosensing and diagnostic applications, a key objective is to design detection systems capable of identifying targets at very low concentrations, i.e., achieving high sensitivity. Here, we propose a linker-mediated detection scheme in which the presence of multivalent target molecules (linkers) facilitates the adsorption of ligand-coated guest nanoparticles onto a receptor-coated host substrate. Through a combination of computer simulations and mean-field theory, we demonstrate that, at fixed overall binding strength, increasing the valency of linkers exponentially lowers the concentration threshold for detection. This counterintuitive behavior arises from the combinatorial entropy associated with multivalent binding configurations, which tremendously amplifies the adsorption sensitivity and enables the identification of targets at extremely low concentrations. Our findings highlight multivalency engineering of linkers as a powerful strategy to substantially enhance the sensitivity of biodetection systems.
multivalent linkers $|$ ultra-sensitive detection $|$ superselectivity
Introduction
Biomarker and pathogen detections are pivotal for early-stage biomarker and pathogen diagnostics Shaw et al. (2007); Jain and Jain (2021). Yet conventional platforms such as ELISA and PCR often require labor-intensive sample preparation and specialized infrastructure, which hampers rapid, low-cost deployment in complex sample matrices Engvall and Perlmann (1971); Mullis and Faloona (1987); Aydin (2015); Bustin and Nolan (2004). A central challenge is to obtain ultrasensitive, superselective, and fast readouts at physiologically low analyte concentrations Wulfkuhle et al. (2003). Multivalency offers an important route to this goal: by forming multiple, individually weak bonds, a probe achieves strong adsorption and sharp, nonlinear responses to target abundance Kitov and Bundle (2003); Huskens (2006); Martinez-Veracoechea and Frenkel (2011). Beyond the classical picture of multivalent guests adsorbing on receptor-coated host substrates Martinez-Veracoechea and Frenkel (2011), many experimentally relevant assays use linker-mediated architectures, in which soluble target molecules (linkers) bridge a ligand-coated guest to a receptor-decorated host. Prominent examples include genome detection using DNA-functionalized colloids whose sticky ends and mobile linkers enable “sandwich” binding, antibodies that bridge antigens and Fc-family receptors, e.g., via engineered bispecific or multivalent formats, and glyco-functionalized nanoparticles engaging lectin receptors on cell membranes Nimmerjahn and Ravetch (2008); Labrijn et al. (2019); Brinkmann and Kontermann (2017); Kiessling and Grim (2013); Wrapp et al. (2020); Budhadev et al. (2020); Curk et al. (2022); Xu et al. (2023); Xia and Ni (2024). However, the design principle of multivalent linkers remains unknown for biosensing. In particular, how does distributing a fixed total binding strength over more linker arms reshape the detection threshold? Here, building on a Langmuir-like adsorption framework Martinez-Veracoechea and Frenkel (2011) and the recent statistical-mechanical theory for multivalent bridging Xia et al. (2025), we develop a linker-mediated detection scheme that achieves both ultrasensitivity and superselectivity.
Model
We consider a coarse-grained model where guest particles adsorb onto a host substrate via multivalent linkers whose two distinct ends bind to ligands on the guest particles and receptors on the host substrate, respectively (Fig. 1a). Following a Langmuir-like framework Martinez-Veracoechea and Frenkel (2011), we consider a substrate site with $n_{r}$ immobile receptors that can accommodate at most one guest particle. Thus, each site exists either in a desorbed state, where no linker bridge is formed, or in an adsorbed state, where at least one linker bridges between the guest particle and receptors on the site. A guest particle is coated with $n_{l}$ mobile ligands, where the mobility of ligands allows them to potentially bridge with any receptor at its corresponding site Xia and Ni (2024). Linkers are treated as an ideal gas of density $\rho=e^{\beta\mu}/\Lambda^{3}$, with the chemical potential $\mu$, de Broglie wavelength $\Lambda$, and the inverse thermal energy $\beta=1/k_{B}T$, where $k_{B}$ and $T$ are the Boltzmann constant and system temperature, respectively. Each linker possesses $\kappa_{l}$ and $\kappa_{r}$ binding ends for ligands and the receptors, with individual binding free energies $f_{l}$ and $f_{r}$, respectively. The formation of a bridging linker is penalized by a mean-field configurational free energy cost $f_{\rm cnf}$.
We formulate the grand potential of the system in both adsorbed and desorbed states. Our approach is based on the framework in Ref. Xia et al. (2025), which has been validated by thermodynamic integration for systems with explicit multivalent linkers. The central idea is to treat each linker as interacting only with receptors on a single site and ligands on the associated guest particle, assuming that the linker size is small compared to the site and guest particles. Thus, linkers mediate an effective interaction between the host and the guest, without coupling with other sites. In the adsorbed state, besides free linkers in solution, each linker may dangle from a receptor, dangle from a ligand, or bridge between receptors and ligands, thereby connecting the substrate and the particle. By applying the saddle-point approximation, we determine the equilibrium number of linkers in each state: $M_{l}$ for those dangling on ligands, $M_{r}$ for those dangling on receptors, and $Q$ for those that form bridges. The resulting grand potential for the adsorbed state is Supplementary Materials
$\beta F_{\rm ad}=\sum_{i\in\{l,r\}}\left(n_{i}\log\bar{n}_{i}-\bar{n}_{i}-M_{i}\right)-Q,$ (1)
where $\bar{n}_{i}$ denotes the number of unbound ligands ($i=l$) or receptors ($i=r$) in the adsorbed state. Similarly, for the desorbed state, only dangling linkers exist, as no bridges can form when the guest particle is absent. The grand-canonical potential in the desorbed state is given by
$\beta F_{\rm de}=\sum_{i\in\{l,r\}}\left(n_{i}\log\bar{n}_{i}^{\prime}-\bar{n}_{i}^{\prime}-M_{i}^{\prime}\right),$ (2)
where $M_{l}^{\prime}$ and $M_{r}^{\prime}$ represent the numbers of linkers dangling from ligands and receptors, respectively, and $\bar{n}_{i}^{\prime}$ represents the number of unbound ligands or receptors when the guest particle is infinitely far from the site. $M$, $Q$, and $\bar{n}$ are numerically calculated by solving a set of self-consistent equations, in which the convergence of the solution for arbitrary linker valency is ensured Xia et al. (2025).
We assume that adsorption events on individual substrate sites are independent, with no interaction between particles adsorbed on adjacent sites. For each site, we take the desorbed state as the reference, and define the referenced partition function of the adsorbed state as $q=e^{-\beta(F_{\rm ad}-F_{\rm de})}-1$ Angioletti-Uberti (2017). The averaged grand partition function of all sites is then given by $\Xi=\langle 1+z_{g}q\rangle_{\langle n_{r}\rangle}$, where $z_{g}$ is the activity of guest particles in the bulk solution, proportional to the particle density and adsorption volume Martinez-Veracoechea and Frenkel (2011); Xia and Ni (2024). $\langle\cdot\rangle_{\langle n_{r}\rangle}$ calculates the average over the Poisson distribution of uncorrelated, immobile receptors, characterized by the mean $\langle n_{r}\rangle$. The probability of nanoparticle adsorption follows a Langmuir-type isotherm:
$\theta:=\frac{\partial\log\Xi}{\partial\log z_{g}}=\left\langle\frac{z_{g}q}{1+z_{g}q}\right\rangle_{\langle n_{r}\rangle}.$ (3)
To validate our theoretical prediction and provide a parameter mapping for future studies, we performed Grand Canonical Monte Carlo (GCMC) simulations of guest nanoparticles in contact with the host substrate coated with explicit receptors. Due to the low density and the multivalent nature of guest particles in most experiments, the equilibration time for 3D simulations is extremely long. Therefore, we simulate the a monolayer of guest particles on the substrate where the guest particles can be modelled as 2D hard disks, and they are in chemical equilibrium with a bulk 3D reservoir above. The adsorption and desorption of guest particles are treated as the insertion and deletion of particles from the reservoir in the simulation Supplementary Materials . Guest particles are modeled as hard disks with diameter $\sigma=1$ on a two-dimensional substrate in chemical equilibrium with a reservoir at chemical potential $\mu_{g}$ Xia et al. (2023); Xia and Ni (2024) (see Fig. 1b). The activity of guest particles is given by $z_{g}=e^{\beta\mu_{g}}\pi\sigma^{2}h_{0}/(4\Lambda^{3})$, where $h_{0}=1$ is the thickness of the adsorption layer. Immobile receptors are randomly distributed on the substrate following a Poisson distribution with number density $\rho_{r}$. Mobile ligands on each guest particle can interact with receptors located within the area underneath the particle. Guest particles move on a two-dimensional plane of box length $L$ with periodic boundary conditions and interact via hard-disk repulsion. The maximum number of adsorption sites is set to $N_{\rm max}=4L^{2}/(\pi\sigma^{2})$, and the adsorption probability is calculated as $\theta=\langle N_{g}\rangle/N_{\rm max}$, where $\langle N_{g}\rangle$ is the average number of adsorbed guest particles.
Figure 1: Multivalent linker-mediated adsorption of guest particles. (a) Schematic illustration of the multivalent linker-mediated interaction between guest nanoparticles and the host substrate. Guest particles that form bridges to the substrate via linkers are considered in the adsorbed state, while those without such bridges are in the desorbed state. (b) Schematic of the GCMC simulation setup. Black dots represent immobile receptors randomly distributed on the substrate. Large gray circles indicate guest particles, each of which can interact only with receptors located within its own shaded region via linkers.
We focus on how linker multivalency modulates two key features of the adsorption systems: sensitivity and selectivity. Here, sensitivity refers to the lowest detectable value of a given environmental parameter $\chi^{\ast}$ at which adsorption becomes significant. Operationally, we define the sensitivity threshold $\chi^{\ast}$ by the condition $\theta(\chi^{\ast})=0.1$. Selectivity quantifies the degree, to which adsorption responds to changes in $\chi$, and is characterized by the selectivity parameter $\alpha_{\chi}={\mathrm{d}\log\theta}/{\mathrm{d}\log\chi}$. When $\alpha_{\chi}>1$, the adsorption increases superlinearly with respect to $\chi$ with $\theta\sim\chi^{\alpha_{\chi}}$. Conversely, $\alpha_{\chi}<-1$ indicates that the desorption also proceeds in a superlinear fashion. Both situations are referred to as $\chi$-dependent superselectivity.
Figure 2: $\rho$-dependent superselectivity and ultra-sensitivity. Here, $\kappa\beta f_{l}=\kappa\beta f_{r}=-5$, $n_{l}=\langle n_{r}\rangle=10$, $\beta f_{\rm cnf}=2$, and $z_{g}=10^{-5}$. The symbols are simulation results, and the solid curves are the theoretical prediction (Eq. 3). (a) The adsorption probability $\theta$ as a function of $\mathrm{log}\rho$ under different valencies of linkers. The dashed lines indicate the low-concentration and high-concentration limits of $\theta$ Supplementary Materials . (b) The selectivity $\alpha_{\rho}$ to linker concentration as a function of $\mathrm{log}\rho$ under different valencies from theoretical prediction Supplementary Materials . (c) The maximal and minimal selectivity $(\alpha_{\rho})_{\mathrm{max/min}}$ as a function of linker valency $\kappa$ from theoretical prediction. The dashed lines are guides of eyes where the curves converge to (d) Log of the adsorption and desorption $\mathrm{log}\rho_{\rm ad,dp}$ as a function of linker valency $\kappa$. The small triangle shows the converging slope of the desorption curve (Eq. 5).
Ultrasensitive $\rho$-response
To investigate how linker valency $\kappa$ affects the adsorption of guest nanoparticles, we perform both theoretical analysis and GCMC simulations, considering symmetric linkers with $\kappa=\kappa_{l}=\kappa_{r}$. We fix the overall binding strength to isolate the effect of linker valency so that any change in the adsorption of nanparticles arises solely from the combinatorial entropy associated with multivalency. Accordingly, in all calculations we fix the total binding strength as $f_{\rm tot}=\kappa f_{l}=\kappa f_{r}$.
As shown in Fig. 2a, the adsorption $\theta$ increases linearly with $\rho$ at low concentrations Supplementary Materials , followed by a pronounced superlinear regime. Remarkably, at high linker concentrations, after a saturation plateau, $\theta$ begins to decrease as $\rho$ increases further, which is a distinct feature of linker-mediated interactionsXia et al. (2020); Lowensohn et al. (2019). The corresponding selectivity parameter $\alpha_{\rho}$, shown in Fig. 2b, reveals that for small and moderate linker valency $\kappa$, both adsorption and desorption transitions are superselective, i.e., the maxima and minima of $\alpha_{\rho}$ exceed 1 in magnitude ($\alpha_{\rho,\max}>1$, $\alpha_{\rho,\min}<-1$). Interestingly, as $\kappa$ increases, the selectivity weakens: $\alpha_{\rho,\max}$ decreases and $\alpha_{\rho,\min}$ increases, both approaching $1$ and 1-1 in the large valency limit (Fig.2c). This is because, as the linker valency increases, fewer linker molecules are required to bridge, and eventually a single linker may suffice to fully bridge the host and guest, so that the superselectivity with respect to linker concentration vanishes. In the limit of zero particle activity $z_{g}\to 0$ and strong individual binding $f_{l/r}\to$-$\infty$, our mean-field analysis predicts that the maximal selectivity decays as $\alpha_{\rho,\max}\approx\max[(\min(n_{l},n_{r}))/\kappa$,1] Supplementary Materials , indicating that the number of bridges ultimately limits the attainable superselectivity. When $\kappa$ exceeds $\min(n_{l},n_{r})$, the system loses its superselective response to linker concentration. In realistic regimes of finite $f_{l/r}$ and $z_{g}$, the selectivity is reduced even further, consistent with MC simulations.
In addition to selectivity, linker valency also strongly influences the adsorption and desorption sensitivity thresholds, $\rho_{\rm ad}$ and $\rho_{\rm dp}$. As shown in Fig. 2d, $\rho_{\rm ad}$ decreases exponentially with increasing $\kappa$, whereas $\rho_{\rm dp}$ decreases and then reach a plateau at large $\kappa$. Since the overall binding strength is fixed, these trends arise from the growing combinatorial entropy introduced by linker multivalency. To elucidate the underlying physics, we examine the large-$\kappa$ limit at a single site.
At the adsorption sensitivity threshold $\rho_{\rm ad}$, both adsorbed and bridging linkers are extremely sparse, because only a very small number of linkers are sufficient to fully bind the guest particle to the site. Thus, the effective interaction between a guest particle and its site mainly depends on the number of unbound ligands and receptors Supplementary Materials
$\beta F_{\rm eff}\simeq\sum_{i=l,r}\left[n_{i}\log\left(\frac{\bar{n}_{i}}{n_{i}}\right)-(\bar{n}_{i}-n_{i})\right].$ (4)
Under the Poisson distribution of receptor numbers, this yields the adsorption sensitivity threshold Supplementary Materials
$\log\rho_{\rm ad}\sim-\lim_{\kappa\to+\infty}\left[\mathrm{log}\left(\bar{n}_{l}+1\right)+\mathrm{log}\left(\bar{n}_{r}+1\right)\right]\kappa.$ (5)
This shows that $\rho_{\rm ad}$ decreases exponentially with increasing linker valency $\kappa$, with the exponent given by the constant $\left[\log(\bar{n}_{l}+1)+\log(\bar{n}_{r}+1)\right]$ at the large $\kappa$ limit, which agrees quantitatively with GCMC simulations (Fig. 2d). Thus, as the total binding strength $\kappa f_{l/r}$ is fixed, this exponential scaling reflects the dominant contribution of the combinatorial entropy associated with linker multivalency. This intriguing result suggests that even with a fixed overall binding strength, increasing $\kappa$ can exponentially lower the adsorption threshold or detection limit, thereby enabling ultrasensitive detection.
Different from the effect of valency $\kappa$ on the adsorption of nanoparticles, the desorption threshold $\rho_{\rm de}$ decreases with increasing $\kappa$ and reaches a plateau at large $\kappa$ limit (Fig. 2d). This is because all receptors and ligands are fully occupied at large $\kappa$. Further increasing $\kappa$ does not increase the number of bound ends per linker, thus has no effect on the adsorption probability $\theta$ and corresponding $\rho_{\rm de}$ Supplementary Materials . This behavior is qualitatively consistent with the purely entropy-driven phase behavior observed in programmable colloidal atom-electron equivalents at large linker valency Xia et al. (2025).
Detection among non-specific binders
Figure 3: Influence of non-specific binders. Here, $\kappa\beta f_{l}=\kappa\beta f_{r}=-5$, $\kappa=3$, $\beta f_{m}=1$, $\beta f_{\rm cnf}=2$, and $z_{g}=10^{-5}$. The symbols are simulation results, while the solid curves are the theoretical prediction (Eq. 3). (a) The adsorption probability $\theta$ as a function of $\log\rho$ for different non-specific binder chemical potential $\mu_{m}$, where $n_{l}=\langle n_{r}\rangle=10$. (b) The adsorption probability $\theta$ as a function of $\log\rho$ under different average receptor numbers $n_{r}$ with $n_{l}=10$ and $\beta\mu_{m}=3$. (c) The adsorption probability $\theta$ as a function of $\log\rho$ under different nanoparticle valencies $n_{l}$ with $\langle n_{r}\rangle=10$ and $\beta\mu_{m}=3$.
In complex environments such as biological fluids or environmental samples, numerous molecules other than the targeted linkers are present Shrivastav et al. (2021); Sampad et al. (2024); Sangkham (2021). These untargeted species can adsorb onto sensor surfaces or recognition elements with weak non-specific interactions. Generally, such non-specific binding would introduce a background signal and reduce the selectivity and sensitivity of the detection system. Hence, careful consideration and mitigation of the environmental non-specific binding is essential for developing a reliable detection system in real-world applications. To this end, we generalize our model by considering non-specific binders. Those non-specific species are modeled as weak monovalent binders that can bond weakly with either ligands or receptors but cannot form bridges between them Fenzl et al. (2016). Here we only consider the non-specific binders, as they are the most common environmental noise in bio-detections, and the untargeted molecules bridging between ligands and receptors are rare Algar et al. (2019). Those binders are present in the environment with individual non-specific binding free energies $f_{m}$ and treated as an ideal gas of density $\rho_{m}=e^{\beta\mu_{m}}/\Lambda^{3}$ with the chemical potential $\mu_{m}$.
We formulate the grand potential of the system with the mixture of the target linkers and non-specific binders by applying the saddle-point approximation. The resulting grand potential for the adsorbed state is Supplementary Materials
$\beta F_{\rm ad}=\sum_{i\in\{l,r\}}\left(n_{i}\log\bar{n}_{i}-\bar{n}_{i}-M_{i}-B_{i}\right)-Q,$ (6)
where $B_{i}$ is the number of adsorbed binders onto the substrate $(i=r)$ or the particle $(i=l)$. Likewise, for the desorbed state, where no bridge forms, the grand potential is
$\beta F_{\rm dp}=\sum_{i\in\{l,r\}}\left(n_{i}\log\bar{n}_{i}^{\prime}-\bar{n}_{i}^{\prime}-M_{i}^{\prime}-B_{i}^{\prime}\right),$ (7)
where $B_{i}^{\prime}$ is the number of adsorbed binders when the particle is far from the substrate.
The theoretical predictions and the simulation results for different chemical potential of binders are shown in Fig. 3a. One can see that with increasing the chemical potential of the non-specific binders, the adsorption of guest nanoparticles mediated by targeted linkers is significantly disrupted, and the saturation range gradually disappears with the detection threshold $\rho_{\rm ad}$ moving to a larger value, which indicates the vanishing of the ultra-sensitivity. This suppression of linker-mediated adsorption is due to the competition between non-specific binders and targeted linkers for binding sites. By adsorbing onto ligands and receptors, the non-specific binders effectively reduce the number of available binding partners for the bridge formation. Therefore, a strategy to mitigate this effect is to increase the total number of ligands or receptors, thereby restoring available binding sites. As shown in Fig. 3b and c, increasing the densities of receptors grafted on host substrate or the number of ligands coated on guest nanoparticles progressively recovers the detection sensitivity of targeted linkers.
Conclusion and Discussions
In summary, we have developed a theoretical framework that elucidates how linker multivalency governs the sensitivity and selectivity of bio-detection systems. Our analysis shows that, at a fixed overall binding strength, increasing linker valency exponentially lowers the detection threshold, thereby enabling ultrasensitive readouts even at extremely low target concentrations. As a result, even sparse populations of multivalent targets can generate robust adsorption signals, which is highly desirable for early-stage biomarker and pathogen diagnostics. At the same time, increasing linker valency reduces superselectivity with respect to linker concentration, reflecting a fundamental trade-off between sensitivity and concentration-based discrimination. Furthermore, the presence of non-specific binders can suppress ultrasensitivity when their concentration becomes sufficiently high. This effect can be mitigated by increasing the densities of receptors grafted on host substrate or the number of ligands coated on guest nanoparticles, thereby restoring the detection sensitivity. Taken together, these findings indicate that optimal detection performance does not necessarily require ligands and receptors with the strongest possible binding affinities. Instead, those capable of forming multiple simultaneous interactions with the target may be more effective. This principle also rationalizes the recently developed ultrasensitive whole-genome detection strategy based on DNA-functionalized colloids, where long single-stranded DNAs (ssDNAs) act as “multivalent linkers”, driving colloidal aggregation at extremely low ssDNA concentrations Xu et al. (2023). Our results provide general design principles for engineering multivalent detection systems and highlight the fundamental role of combinatorial entropy in biomolecular recognition. The theoretical insights presented here offer practical guidelines for the development of next-generation biosensors and diagnostic assays capable of ultra-sensitive and highly specific detection in complex biological environments.
Acknowledgements.
X. X. acknowledges support from the Alexander von Humboldt-Stiftung. This work was financially supported by the Academic Research Fund from the Singapore Ministry of Education (RG151/23, and RG88/25) and the National Research Foundation, Singapore, under its 29th Competitive Research Program (CRP) Call (NRF-CRP29-2022-0002).
I Data Availability Statements
The data that support the findings of this article are openly available https://doi.org/10.5281/zenodo. (17426124).